HELP! Any Math Wiz out there? Fractions get me everytime!
solve for t showing your work.
2/3+1/t=4/5

Found 2 solutions by Alan3354, edjones:
Answer by Alan3354(69443)

2/3+1/t=4/5
Multiply thru by 15t
10t + 15 = 12t
2t = 15
t = 7.5
---------
Or,
1/t = (4/5) - 2/3) = 12/15 - 10/15
1/t = 2/15
t = 15/2

Answer by edjones(8007)

2/3+1/t=4/5
10t+15=12t
-2t=-15
t=15/2
.
Ed